A 44.0-cm diameter disk rotates with a constant angular acceleration of 2.60 rad/s2. It starts from rest at t = 0, and a line drawn from the center of the disk to a point P on the rim of the disk makes an angle of 57.3° with the positive x-axis at this time.


(a) At t = 2.47 s, find the angular speed of the wheel.

 __rad/s


(b) At t = 2.47 s, find the magnitude of the linear velocity and tangential acceleration of P.

linear velocity ___ m/s

tangential acceleration ___ m/s2


(c) At t = 2.47 s, find the position of P (in degrees, with respect to the positive x-axis).

___ ° counterclockwise from the +x-axis




1
Expert's answer
2021-10-05T10:07:49-0400

Radius of the disk r=0.22m

Angular acceleration "\u03b1=2.6rad\/s^2"

Initial angle "\u03b8_i=57.3^\u00b0"

Time difference Δ=2.47s

a) The angular speed of the wheel is

"\u03c9_f =\u03c9_i + \u03b1t"

Here, "\u03c9_i=0" since it starts from rest

"\u03c9_f =0rad\/s +(2.6rad\/s^2)(2.47s)= 6.422 rad\/s"

b) The linear velocity of the point P is

"v=rω_f \\

=(0.22m)(6.422rad/s) \\

=1.4128m/s"

The tangential acceleration is

"a_t=r α \\

=(0.22m)(2.6rad/s^2) \\

=0.572m/s^2"

c) The position of the point after 2.47s is

"Δθ= ω_it +\frac{1}{2}αt^2 \\

=0+\frac{1}{2}(2.6rad/s^2)(2.47s)^2 \\

=7.931rad"

In degrees,

"θ_f=θ_i +(7.931rad)(\frac{180^°}{π} rad) \\

=57.3^° + 455^° \\

=513.3^°"

From x-axis it would be

513.3 – 360= 153.3

The angle is "153.3^\u00b0" from x-axis